# Information geometry approach to quantum stochastic thermodynamics

### v3

@@ I Introduction
Quantum systems often require a more refined approach compared to their classical counterparts due to quantum correlations that are captured by density matrices rather than classical probability distributions. This has interesting consequences in information geometry: while the metric in classical probability spaces is unique [32], the states of quantum systems live within a Hilbert space where a variety of metrics can be used to measure distances between states, depending on the chosen inner product. Petz’s work on monotone metrics has been significant in this regard. He investigated the conditions under which metrics on the space of density matrices remain monotone under completely positive, trace-preserving (CPTP) maps [42, 43, 44, 45]. Petz identified a family of such monotone metrics, all of which are contractive under CPTP maps, ensuring that the statistical distance between quantum states does not increase through physical operations, consistent with the principle that information cannot be enhanced merely by processing data [46]. Notable examples include the symmetric logarithmic derivative (SLD) metric (smallest), the Kubo-Mori-Bogoliubov (KMB) metric, the harmonic mean (HM) metric (largest), and the Wigner-Yanase (WY) metric. For a comprehensive review on this topic, we refer readers to [47].
@@
The aim of our work is to extend classical results connecting information geometry and stochastic thermodynamics to quantum systems. To this end, we first show that the uncertainty relation connecting the geometric uncertainty in the path in state space and the time-averaged rate of information change, previously established for classical systems [35], carries over seamlessly to the quantum domain. In fact, we find that the bound can be tightend both in classical and quantum settings. Provided the dynamics is arbitrary differentiable and trace-preserving, any QFI can be split into a metric-independent incoherent and a metric-dependent coherent contribution [48, 49]. Building on this result, we then show that any QFI can be uniquely linked to the entropic acceleration through their shared incoherent component. This allows us to demonstrate that when the quantum system’s dynamics are governed by a GKSL master equation [50, 51], the incoherent part of the QFI can be expressed in terms of both the rate of change in generalized thermodynamic forces, driving probability currents between instantaneous eigenstates of the system’s density matrix, as well as the entropic flux exchanged with an environment, mirroring the established relationship in classical systems [34]. Further, we generalise a classical geometrical bound on the rate of change of the von Neumann entropy to the quantum case by the addition of a non-negative term arising from the coherent contributions to the geometric action. To further illustrate these findings, we analyze the recently proposed quantum thermodynamic Mpemba effect [52] through the lens of information geometry.
@@
The paper is structured as follows: section II reviews the Fisher information with respect to time, emphasizing its geometric interpretation. Section III extends this framework to quantum systems, introducing the family of quantum Fisher informations (QFIs) and their explicit forms for arbitrary differentiable trace-preserving quantum dynamics (section IV). In section V.0.1, we examine classical results linking the Fisher information on the time parameter to entropic acceleration, entropic flow, and entropy production rates. Section V.0.2 presents our first main result, showing how these classical results extend to quantum systems. We then present our second key result, which builds on the first by extending a classical bound on the entropy rate — expressed in terms of geometric quantities describing the path traced by the system’s evolution in state space — by including a non-negative contribution from the coherent part of the geometric action. Finally, in section VII, we illustrate the relationship between quantum thermodynamics and information geometry using the recently reported quantum-thermodynamic Mpemba effect [52]. We demonstrate that the relaxation speed-up is captured by the “ratio of completion”, a quantity grounded in the geometric statistical distance, and analyze the geometric uncertainty. Lastly, we examine a previously derived geometric bound on the time evolution of arbitrary observables [38, 40] with a focus on the Hamiltonian, establishing a connection to the dissipated heat.
@@ II Fisher information with respect to time
$F(t)=\sum_{ij}\frac{\mathrm{d}\theta_{i}}{\mathrm{d}t}m_{ij}\frac{\mathrm{d}\theta_{i}}{\mathrm{d}t}=\sum_{x}p_{x}(t)\left[\frac{\mathrm{d}\log p_{x}(t)}{\mathrm{dt}}\right]^{2}.$ (2)
@@ III Quantum extensions
In quantum settings, the objects of interest are states represented by density matrices rather than probability distributions. Like in the classical case, a statistical line element with respect to time $t$ in quantum state space may be defined, however, the choice of the metric is not unique. Rather, there exists a family of metrics, all interpretable as a different QFI, characterised by the Morozova, Čencov and Petz theorem [43, 44, 42, 58, 59, 47]. According to this characterization, any metric contractive under evolution must yield a squared line element of the form (up to multiplication by a scalar constant) [43]
@@ IV General time evolution
The spectral decomposition of the state at time $t$ and the initial state are, respectively, $\hat{\rho}_{t}=\sum_{x}p_{x}(t)\ket{x(t)}\bra{x(t)}$ | and $\hat{\rho}_{0}=\sum_{x}p_{x}(0)\ket{x(0)}\bra{x(0)}$ |. The unitary operator $U_{t}$ transforms between the respective eigenbasis elements, so that |$\ket{x(t)}=\hat{U}_{t}\ket{x(0)}$. The matrix $\hat{\chi}_{t}$ is given by $\hat{\chi}_{t}=\sum_{x}p_{x}(t)\ket{x(0)}\bra{x(0)}$ |, i.e. it is the matrix that is diagonal in the original basis, but with the time evolved spectrum on the diagonal. Eq. 11 simply expresses that the state might have a time-evolving spectrum (through incoherent processes) in a time-evolving eigenbasis (through coherent evolution). This allows us to define the Hermitian operator $\hat{H}_{t}=i(\mathrm{d}\hat{U}_{t}/\mathrm{d}t)\hat{U}_{t}^{\dagger}$, which may be viewed as an effective “Hamiltonian”, and the differential equation governing the evolution of $\hat{\rho}_{t}$ is [62, 56, 63]
@@
Coming back to the QFI, we note that the factor common to every member of the QFI family is the change of the matrix element $|\partial_{t}\hat{\rho}_{xy}|^{2}$ in the instantaneous eigenbasis of $\hat{\rho}_{t}$. Let us plug in the form in Eq. 12:
@@
Interestingly, the coherent contribution of any of the QFI variants is a measure of coherence with respect to the time-dependent Hermitian operator $\hat{H}_{t}$ [64, 36].
Since the square infinitesimal length element is given by $\mathrm{d}s^{2}=F_{Q}(\hat{\rho}_{t})\mathrm{d}t^{2}$, the statistical distance in state space is given by
@@ V.0.1 Classical systems
where $w^{xy}(t)>0$ (if $x\neq y$) is the transition rate from state $y\to x$, and $w^{yy}(t)=-\sum_{x\neq y}w^{xy}(t)$, which leads to $\sum_{x}w^{xy}=0$, ensuring the normalisation of $p_{t}$. This allows us to rewrite the master equation in terms of the probability currents from state $y$ to $x$, $j^{xy}(t)=w^{xy}(t)p_{y}(t)-w^{yx}(t)p_{x}(t)$, as
@@
$\Phi(t)$ $=\left\langle\left\langle\phi\right\rangle\right\rangle.$ (30)
@@
We now take a closer look at the term $\mathcal{B}=-\sum_{x}\ddot{p}_{x}(t)\log p_{x}(t)$: to this end, we first note that the the rate of change of the Shannon entropy $\dot{S}(t)=-\langle\langle I(t)\rangle\rangle$ can be thought of as the rate of change of the average information. Therefore, the Shannon entropy’s second derivative $\ddot{S}(t)=-\frac{\mathrm{d}}{\mathrm{d}t}\langle\langle I(t)\rangle\rangle$ in turn tells us how the rate of change of the average information changes in time — it is the rate of change of the current averaged relative surprisal. The Fisher information, however, is an average over the rate of change in information between each set of states $x$ and $y$, that is $F(t)=\langle\langle\dot{I}(t)\rangle\rangle$ — it is the current average over the rate of change of the relative surprisal. Therefore, $\mathcal{B}=\left\langle\left\langle\frac{\mathrm{d}}{\mathrm{d}t}I(t)\right\rangle\right\rangle-\frac{\mathrm{d}}{\mathrm{d}t}\left\langle\left\langle I(t)\right\rangle\right\rangle$ is the difference of the average local information rate and the bulk information rate. Using the definition in Eq. 28 one finds [34]
@@ V.0.2 Quantum generalisation
Note here, that the Hamiltonian $\hat{H}_{t}$ generally does not coincide with the Hermitian time-dependent operator introduced in section IV, but rather can be reconstructed from the eigenstates of the time-evolved state $\hat{\rho}$ from which $\hat{U}_{t}$ is obtained.
@@
In section IV we have illustrated that the QFI with respect to the time parameter can always be split into an incoherent and a coherent contribution. Crucially, the incoherent contribution dependents solely on the instantaneous spectrum of the state. We may thus play the same game as in the classical case and relate the incoherent part of the QFI to thermodynamic observables.
Again, we may express $\dot{S}(t)$ in terms of the probability currents and further split it into the entropy production rate $\sigma$ and the entropic flow rate to the environment $\Phi$
@@
where $\mathcal{B}=-\sum_{x}\ddot{p}_{x}\log p_{x}$, mirroring the classical result in [35]. Therefore, it may alternatively expressed as [34]
@@ VI Geometric bound on the von Neumann entropy rate
The above bound constitutes the second key result of our work. Similar to the classical case [35], the change in the entropy rate is constrained by the sum of the integrated difference between the average local information rate and the bulk information rate, as well as the geometric uncertainty. However, an additional non-negative quantum contribution emerges, which we identify as the geometric action associated with the coherent dynamics. As briefly described in Sec. IV, the action is associated to the ‘energy’ of the path due to its similarity with the kinetic energy integral in classical mechanics. It is worth pointing out that the first two terms vanish only in stationary states [35], whereas the last tern vanishes in stationary states or whenever the dynamics is purely incoherent. Therefore, their sum vanishes only in stationary states. Note that if one considers a driving protocol where the initial state is the fixed point of the initial instantaneous generator of the dynamics, such that $\dot{S}(0)=0$, the above serves as a bound on the instantaneous rate of entropy change $\dot{S}(T)$ at any later time $T$, which indicates that the state is out of equilibrium. We can further exploit the relationship between the incoherent contribution to the QFI and the entropic acceleration in Markovian open system dynamics obeying local detailed balance. In cases where the heat flow between the system and its environment vanishes, the resulting bound applies to the entropy production. Conversely, if the entropy production is zero, the bound constrains the entropic flow.
@@ VII Quantum Mpemba effect
$F_{\mathrm{neq}}(\hat{\sigma})=\beta^{-1}\mathrm{D}(\hat{\sigma}\|\hat{\tau}_{\beta})+F_{\mathrm{eq}},$ (49)
with equilibrium free energy $F_{\mathrm{eq}}=-\beta^{-1}\log Z_{\beta}$, partition function $Z_{\beta}=\mathrm{Tr}\left[\exp(-\beta\hat{H})\right]$ and $\mathrm{D}(\hat{\sigma}\|\hat{\tau}_{\beta})$ is the quantum relative entropy between the relaxing state and the thermal fixed point $\hat{\tau}_{\beta}=\exp(-\beta\hat{H})/Z_{\beta}$, with respect to the system Hamiltonian $\hat{H}$. The authors in [52] defined a Mpemba effect to occur when a $\hat{\rho}_{0}^{\prime}$ with higher initial non-equilibrium free energy $F_{\mathrm{neq}}$ compared to the reference state $\hat{\rho}_{0}$, relaxes to thermal equilibrium faster. In particular, there exists a time $t_{M}$ after which $F_{\mathrm{neq}}(\hat{\rho}(t))>F_{\mathrm{neq}}(\hat{\rho}^{\prime}(t))$ for all $t>t_{M}$, indicating a crossing of non-equilibrium free energies along their respective relaxation paths.
Figure 1: a) Time-evolution of the total SLD QFI of the reference state $\hat{\rho}$ (blue) and the rotated state $\hat{\rho}^{\prime}$ (red). b) The rotated state evolves classically (red), therefore the quantum Fisher information reduces to the incoherent contribution. The reference state, however, is coherent in the energy eigenbasis and thus has both incoherent (dashed blue) and coherent (dash-dotted). We find that the incoherent contribution decays exponentially faster than the coherent contribution.
@@
Note that, in this context, the interaction with the environment cannot generate coherence in the energy eigenbasis. Thus, under dynamics induced by the GKSL master equation (Eq. 50), the rotated state will remain diagonal in the eigenbasis of the Hamiltonian at all later times. We now remember that a general time evolution can be understood as the combination of time-evolving eigenvalues, represented by $\chi(t)$ in a time evolving eigenbasis, expressed in $U_{t}$ (see section IV). But since the eigenbasis remains static, $U_{t}$ is the identity, and the dynamics are fully captured by the time-evolving eigenvalues. Therefore, the QFI with respect to time reduces to its incoherent contribution - the classical Fisher information for the probability distribution constructed from the time-evolving eigenvalues. The reference state, however, initially has coherence in the energy eigenbasis by construction, and therefore the QFI has both an incoherent and a coherent contribution, as shown in Fig. 1b).
@@
Interestingly, for the SLD geodesic lengths (see Eq. 9) we find the opposite to be true (dashed lines), as shown in Fig. 2a). Although the rotated state does not evolve along a geodesic - meaning the QFI is not constant along the path - the length of its path matches that of a geodesic at each instant, which describes a strictly monotonic trajectory along the $z$-axis of the Bloch sphere. Additionally, we find that the ratio of completion, indicating the fraction of the statistical distance traced as a function of time is higher for the rotated state than for the reference state $\mathcal{R}^{\prime}_{\tau}(t)=\mathcal{L}_{\mathrm{SLD}}(\hat{\rho}^{\prime}(t))/\mathcal{L}_{\mathrm{SLD}}(\hat{\rho}^{\prime}(\tau)>\mathcal{R}_{\tau}(t)=\mathcal{L}_{\mathrm{SLD}}(\hat{\rho}(t))/\mathcal{L}_{\mathrm{SLD}}(\hat{\rho}(\tau))$, as shown in Fig. 2 b), indicating further that the rotated state reaches equilibrium faster than the reference state.
@@
expressing that the time-integrated ratio between the change in the observable expectation value due to state changes, $\left\lvert\dot{o}(t)\right\rvert=\left\lvert\mathrm{Tr}\left[\hat{O}\frac{\mathrm{d}\hat{\rho}_{t}}{\mathrm{d}t}\right]\right\rvert$, and the standard deviation of the observable with respect to the instantaneous state $\hat{\rho}_{t}$, $\Delta_{\hat{\rho}_{t}}\hat{O}=\sqrt{\mathrm{Tr}\left[\hat{\rho}_{t}\hat{O}^{2}\right]-\mathrm{Tr}\left[\hat{\rho}_{t}\hat{O}\right]^{2}}$, is upper bounded by twice the SLD statistical distance $\mathcal{L}_{\mathrm{SLD}}$. Note that since $\mathcal{L}_{\mathrm{SLD}}\geq\mathcal{L}_{\mathrm{SLD}}^{\mathrm{geo}}$ (see Eq. 9), the ratio integrated over a geodesic path yields the lowest bound.

### v4

@@ I Introduction
Quantum systems often require a more refined approach compared to their classical counterparts due to quantum correlations that are captured by density matrices rather than classical probability distributions. This has interesting consequences in information geometry: while the metric in classical probability spaces is unique [32], the states of quantum systems live within a Hilbert space where a variety of metrics can be used to measure distances between states, depending on the chosen inner product. Petz’s work on monotone metrics has been significant in this regard. He investigated the conditions under which metrics on the space of density matrices remain monotone under completely positive, trace-preserving (CPTP) maps [42, 43, 44, 45]. Petz identified a family of such monotone metrics, all of which are contractive under CPTP maps, ensuring that the statistical distance between quantum states does not increase through physical operations, consistent with the principle that information cannot be increased merely by processing data [46]. Notable examples include the symmetric logarithmic derivative (SLD) metric (smallest), the Kubo-Mori-Bogoliubov (KMB) metric, the harmonic mean (HM) metric (largest), and the Wigner-Yanase (WY) metric. For a comprehensive review on this topic, we refer readers to [47].
@@
The aim of our work is to extend classical results connecting information geometry and stochastic thermodynamics to quantum systems. To this end, we first show that the uncertainty relation connecting the geometric uncertainty in the path in state space and the time-averaged rate of information change, previously established for classical systems [35], carries over seamlessly to the quantum domain. In fact, we find that the bound can be tightend both in classical and quantum settings. Provided the dynamics is arbitrary differentiable and trace-preserving, any QFI can be split into a metric-independent incoherent and a metric-dependent coherent contribution [48, 49]. Building on this result, we then show that any QFI can be uniquely linked to the entropic acceleration through their shared incoherent component. This allows us to demonstrate that when the quantum system’s dynamics are governed by a GKSL master equation [50, 51], the incoherent part of the QFI can be expressed in terms of both the rate of change in generalized thermodynamic forces, driving probability currents between instantaneous eigenstates of the system’s density matrix, as well as the entropic flux exchanged with an environment, mirroring the established relationship in classical systems [34]. Further, we generalise a classical geometrical bound on the rate of change of the von Neumann entropy to the quantum case by the addition of a non-negative term arising from the coherent contribution to the geometric action. To further illustrate these findings, we analyze the recently proposed quantum thermodynamic Mpemba effect [52] through the lens of information geometry.
@@
The paper is structured as follows: section II reviews the Fisher information with respect to time, emphasizing its geometric interpretation. Section III extends this framework to quantum systems, introducing the family of quantum Fisher informations (QFIs) and their explicit forms for arbitrary differentiable trace-preserving quantum dynamics (section IV). In section V.0.1, we examine classical results linking the Fisher information with respect to the time parameter to entropic acceleration, entropic flow, and entropy production rates. Section V.0.2 presents our first main result, showing how these classical results extend to quantum systems. We then present our second key result, which builds on the first by extending a classical bound on the entropy rate — expressed in terms of geometric quantities describing the path traced by the system’s evolution in state space — by including a non-negative contribution from the coherent part of the geometric action. Finally, in section VII, we illustrate the relationship between quantum thermodynamics and information geometry using the recently reported quantum-thermodynamic Mpemba effect [52]. We demonstrate that the relaxation speed-up is captured by the “ratio of completion”, a quantity grounded in the geometric statistical distance, and analyze the geometric uncertainty. Lastly, we examine a previously derived geometric bound on the time evolution of arbitrary observables [38, 40] with a focus on the Hamiltonian, establishing a connection to the dissipated heat.
@@ II Fisher information with respect to time
$F(t)=\sum_{ij}\frac{\mathrm{d}\theta_{i}}{\mathrm{d}t}m_{ij}\frac{\mathrm{d}\theta_{j}}{\mathrm{d}t}=\sum_{x}p_{x}(t)\left[\frac{\mathrm{d}\log p_{x}(t)}{\mathrm{dt}}\right]^{2}.$ (2)
@@ III Quantum extensions
In quantum settings, the objects of interest are states represented by density matrices rather than probability distributions. Like in the classical case, a statistical line element with respect to time $t$ in quantum state space may be defined, however, the choice of the metric is not unique. Rather, there exists a family of metrics, all interpretable as a different QFI, characterised by the Morozova, Čencov and Petz theorem [43, 44, 42, 58, 59, 47]. According to this characterization, any metric contractive under stochastic evolution must yield a squared line element of the form (up to multiplication by a scalar constant) [43]
@@ IV General time evolution
The spectral decomposition of the state at time $t$ and the initial state are, respectively, $\hat{\rho}_{t}=\sum_{x}p_{x}(t)\ket{x(t)}\bra{x(t)}$ | and $\hat{\rho}_{0}=\sum_{x}p_{x}(0)\ket{x(0)}\bra{x(0)}$ |. The unitary operator $U_{t}$ transforms between the respective eigenbasis elements, so that |$\ket{x(t)}=\hat{U}_{t}\ket{x(0)}$. The matrix $\hat{\chi}_{t}$ is given by $\hat{\chi}_{t}=\sum_{x}p_{x}(t)\ket{x(0)}\bra{x(0)}$ |, i.e. it is the matrix that is diagonal in the original basis, but with the time evolved spectrum on the diagonal. Eq. (11) simply expresses that the state might have a time-evolving spectrum (through incoherent processes) in a time-evolving eigenbasis (through coherent evolution). This allows us to define the Hermitian operator $\hat{H}_{t}=i(\mathrm{d}\hat{U}_{t}/\mathrm{d}t)\hat{U}_{t}^{\dagger}$, which may be viewed as an effective “Hamiltonian”, and the differential equation governing the evolution of $\hat{\rho}_{t}$ is [62, 56, 63]
@@
Coming back to the QFI, we note that the factor common to every member of the QFI family is the change of the matrix element $|\partial_{t}\hat{\rho}_{xy}|^{2}$ in the instantaneous eigenbasis of $\hat{\rho}_{t}$. Let us plug in the form in Eq. (12):
@@
Interestingly, the coherent contribution of any of the QFI variants is a measure of coherence, also referred to as asymmetry, with respect to the time-dependent Hermitian operator $\hat{H}_{t}$ [64, 36].
Since the square infinitesimal length element is given by $\mathrm{d}s^{2}=1/4\,F_{Q}(\hat{\rho}_{t})\mathrm{d}t^{2}$, the statistical distance in state space is given by
@@ V.0.1 Classical systems
where $w^{xy}(t)>0$ (if $x\neq y$) is the transition rate from state $y\to x$, and $w^{yy}(t)=-\sum_{x\neq y}w^{xy}(t)$, which leads to $\sum_{x}w^{xy}(t)=0$, ensuring the normalisation of $p_{t}$. This allows us to rewrite the master equation in terms of the probability currents from state $y$ to $x$, $j^{xy}(t)=w^{xy}(t)p_{y}(t)-w^{yx}(t)p_{x}(t)$, as
@@
$\Phi(t)$ $=\left\langle\left\langle\phi(t)\right\rangle\right\rangle.$ (30)
@@
We now take a closer look at the term $\mathcal{B}=-\sum_{x}\ddot{p}_{x}(t)\log p_{x}(t)$: to this end, we first note that the the rate of change of the Shannon entropy $\dot{S}(t)=-\langle\langle I(t)\rangle\rangle$ can be thought of as the rate of change of the average information. Therefore, the Shannon entropy’s second derivative $\ddot{S}(t)=-\frac{\mathrm{d}}{\mathrm{d}t}\langle\langle I(t)\rangle\rangle$ in turn tells us how the rate of change of the average information changes in time — it is the rate of change of the current averaged relative surprisal. The Fisher information, however, is an average over the rate of change in information between each set of states $x$ and $y$, that is $F(t)=\langle\langle\dot{I}(t)\rangle\rangle$ — it is the current average over the rate of change of the relative surprisal. Therefore, $\mathcal{B}=\left\langle\left\langle\frac{\mathrm{d}}{\mathrm{d}t}I(t)\right\rangle\right\rangle-\frac{\mathrm{d}}{\mathrm{d}t}\left\langle\left\langle I(t)\right\rangle\right\rangle$ is the difference of the average local information rate and the bulk information rate. Using the definition in Eq. (28) one finds [34]
@@ V.0.2 Quantum generalisation
Note here, that the Hamiltonian $\hat{H}_{t}$ generally does not coincide with the Hermitian time-dependent operator introduced in section IV, but rather can be reconstructed from the eigenstates of the time-evolved state $\hat{\rho}_{t}$ from which $\hat{U}_{t}$ is obtained.
@@
In section IV we have illustrated that the QFI with respect to the time parameter can always be split into an incoherent and a coherent contribution. Crucially, the incoherent contribution dependents solely on the instantaneous spectrum of the state. We may thus play the same game as in the classical case and relate the incoherent part of the QFI to thermodynamic observables. Again, we may express $\dot{S}(t)$ in terms of the probability currents and further split it into the entropy production rate $\sigma$ and the entropic flow rate to the environment $\Phi$
@@
where $\mathcal{B}=-\sum_{x}\ddot{p}_{x}\log p_{x}$, mirroring the classical result in [35]. Therefore, it may be alternatively expressed as [34]
@@ VI Geometric bound on the von Neumann entropy rate
The above bound constitutes the second key result of our work. Similar to the classical case [35], the change in the entropy rate is constrained by the sum of the integrated difference between the average local information rate and the bulk information rate, as well as the geometric uncertainty. However, an additional non-negative quantum contribution emerges, which we identify as the geometric action associated with the coherent dynamics. As briefly described in Sec. IV, the action is associated to the ‘energy’ of the path due to its similarity with the kinetic energy integral in classical mechanics. It is worth pointing out that the first two terms vanish only in stationary states [35], whereas the last term vanishes in stationary states or whenever the dynamics is purely incoherent. Therefore, their sum vanishes only in stationary states. Note that if one considers a driving protocol where the initial state is the fixed point of the initial instantaneous generator of the dynamics, such that $\dot{S}(0)=0$, the above serves as a bound on the instantaneous rate of entropy change $\dot{S}(T)$ at any later time $T$, which indicates that the state is out of equilibrium. We can further exploit the relationship between the incoherent contribution to the QFI and the entropic acceleration in Markovian open system dynamics obeying local detailed balance. In cases where the heat flow between the system and its environment vanishes, the resulting bound applies to the entropy production. Conversely, if the entropy production is zero, the bound constrains the entropic flow.
@@ VII Quantum Mpemba effect
$F_{\mathrm{neq}}(\hat{\sigma})=\beta^{-1}D(\hat{\sigma}\|\hat{\tau}_{\beta})+F_{\mathrm{eq}},$ (49)
with equilibrium free energy $F_{\mathrm{eq}}=-\beta^{-1}\log Z_{\beta}$, partition function $Z_{\beta}=\mathrm{Tr}\left[\exp(-\beta\hat{H})\right]$ and $D(\hat{\sigma}\|\hat{\tau}_{\beta})=\mathrm{Tr}\left[\hat{\sigma}\left(\log\hat{\sigma}-\log\hat{\tau}_{\beta}\right)\right]$ is the quantum relative entropy between the relaxing state and the thermal fixed point $\hat{\tau}_{\beta}=\exp(-\beta\hat{H})/Z_{\beta}$, with respect to the system Hamiltonian $\hat{H}$. The authors in [52] defined a Mpemba effect to occur when a $\hat{\rho}_{0}^{\prime}$ with higher initial non-equilibrium free energy $F_{\mathrm{neq}}$ compared to the reference state $\hat{\rho}_{0}$, relaxes to thermal equilibrium faster. In particular, there exists a time $t_{M}$ after which $F_{\mathrm{neq}}(\hat{\rho}(t))>F_{\mathrm{neq}}(\hat{\rho}^{\prime}(t))$ for all $t>t_{M}$, indicating a crossing of non-equilibrium free energies along their respective relaxation paths.
Figure 1: a) Time-evolution of the total SLD QFI of the reference state $\hat{\rho}$ (blue) and the rotated state $\hat{\rho}^{\prime}$ (red). b) The rotated state evolves classically (red), therefore the quantum Fisher information reduces to the incoherent contribution. The reference state, however, is coherent in the energy eigenbasis and thus has both incoherent (dashed blue) and coherent (dash-dotted) contributions. We find that the incoherent contribution decays exponentially faster than the coherent contribution.
@@
Note that, in this context, the interaction with the environment cannot generate coherence in the energy eigenbasis. Thus, under dynamics induced by the GKSL master equation (Eq. (50)), the rotated state will remain diagonal in the eigenbasis of the Hamiltonian at all later times. We now remember that a general time evolution can be understood as the combination of time-evolving eigenvalues, represented by $\chi(t)$ in a time evolving eigenbasis, expressed in $U_{t}$ (see section IV). But since the eigenbasis remains static, $U_{t}$ is the identity, and the dynamics are fully captured by the time-evolving eigenvalues. Therefore, the QFI with respect to time reduces to its incoherent contribution - the classical Fisher information for the probability distribution constructed from the time-evolving eigenvalues. The reference state, however, initially has coherence in the energy eigenbasis by construction, and therefore the QFI has both an incoherent and a coherent contribution, as shown in Fig. 1b).
@@
Interestingly, for the SLD geodesic lengths (see Eq. (9)) we find the opposite to be true (dashed lines), as shown in Fig. 2a). Although the rotated state does not evolve along a geodesic - meaning the QFI is not constant along the path - the length of its path matches that of a geodesic at each instant, which describes a strictly monotonic trajectory along the $z$-axis of the Bloch sphere. Additionally, we find that the ratio of completion, indicating the fraction of the statistical distance traced as a function of time is higher for the rotated state than for the reference state $\mathcal{R}^{\prime}_{\tau}(t)=\mathcal{L}_{\mathrm{SLD}}(\hat{\rho}^{\prime}(t))/\mathcal{L}_{\mathrm{SLD}}(\hat{\rho}^{\prime}(\tau)>\mathcal{R}_{\tau}(t)=\mathcal{L}_{\mathrm{SLD}}(\hat{\rho}(t))/\mathcal{L}_{\mathrm{SLD}}(\hat{\rho}(\tau))$, as shown in Fig. 2 b), indicating further that the rotated state reaches equilibrium faster than the reference state.
@@
expressing that the time-integrated ratio between the change in the observable expectation value due to state changes, $\left\lvert\dot{o}(t)\right\rvert=\left\lvert\mathrm{Tr}\left[\hat{O}\frac{\mathrm{d}\hat{\rho}_{t}}{\mathrm{d}t}\right]\right\rvert$, and the standard deviation of the observable with respect to the instantaneous state $\hat{\rho}_{t}$, $\Delta_{\hat{\rho}_{t}}\hat{O}=\sqrt{\mathrm{Tr}\left[\hat{\rho}_{t}\hat{O}^{2}\right]-\mathrm{Tr}\left[\hat{\rho}_{t}\hat{O}\right]^{2}}$, is upper bounded by twice the SLD statistical distance $\mathcal{L}_{\mathrm{SLD}}$. Note that since $\mathcal{L}_{\mathrm{SLD}}\geq\mathcal{L}_{\mathrm{SLD}}^{\mathrm{geo}}$ (see Eq. (9)), the ratio integrated over a geodesic path yields the lowest bound.